# Streaming Generation of Co-Speech Gestures via Accelerated Rolling Diffusion

### v1

@@ 5.3 User Study
Participants were shown pairs of 15-second videos, each synchronized with the same audio but generated using different models. Both videos were displayed simultaneously, with their positions randomized in each trial to eliminate possible bias towards one side. The participants were asked to compare the two animations and rate them based on style consistency, naturalness and fluidity of animations, audio-animation synchronization, presence of technical issues (such as the gluing effect). The rating could take values $\{-2,-1,0,1,2\}$ where $-2$ indicates a strong preference for the baseline, $-1$ – slight baseline preference, 00 means no noticeable differences, $1$ indicates a slight preference for our model and $2$ indicates a strong preference for our model. To ensure the reliability of our evaluators, we included some videos multiple times and verified that the assessors provided consistent ratings before including their responses in the final analysis. The study was carried out on 60 pairs of video (10 per style in six different styles). Twenty-two professional assessors trained to work with video data were recruited to participate.
@@
The distribution of the user study results is shown in Fig. 2 (Left). Our rolling modification of DSG statistically significantly outperforms original DSG, which correlates with the results of quantitative tests.
5.4 Rolling Diffusion Ladder Acceleration Results
@@
To evaluate the effectiveness of Rolling Diffusion Ladder Acceleration (RDLA) in improving inference efficiency, we conducted a series of experiments on the ZEGGS dataset [9]. This dataset was selected due to its high fidelity, as it is obtained through motion capture, minimizing noise and annotation errors. To ensure a fair comparison, we applied RDLA to the DiffuseStyleGesture (DSG) model [34], which has demonstrated state-of-the-art performance among diffusion-based methods on this benchmark. Following the evaluation protocol in [34], we tested the performance of models on the ZEGGS validation set, which includes six distinct styles: Happy, Sad, Neutral, Old, Angry, and Relaxed. This results in a total of 36 audio samples for evaluation.
@@
Accelerated inference via fewer diffusion steps
A straightforward method for accelerating the sampling process is reducing the number of denoising steps per frame. As described in Sec. 3.5, in our framework, each frame undergoes a predefined number of denoising steps $s$ before the rolling window shifts forward in time. We experimented with reducing this number $s_{r}$ from $s$ to 1 leading to a total inference step count of $T_{r}=s_{r}\cdot N$, where $N$ is the total number of frames in sliding window.
We compared the performance of our accelerated approach against the original DSG method under the same total denoising steps $T_{r}$. Additionally, we evaluated both DDPM [12] and DDIM [30] sampling strategies, as DDIM is known to outperform DDPM when the number of steps gets smaller. Experiments were conducted with the same model settings as in Sec. 5.2 ($N=100$, $n^{\text{cont}}=8$).
@@
As shown in Tab. 3, reducing the number of denoising steps generally maintains and even improves performance in both Fréchet distance (FD) and diversity (Div) metrics, with our approach consistently outperforming the baseline DSG method. Notably, while DDIM achieves competitive results, DDPM remains superior in most cases. These findings suggest that our method can effectively operate with a reduced number of denoising steps (e.g. $T_{r}=100$) without significant degradation in quality.
RDLA experiments
[…]
@@
Results in Tab. 4 indicate that a 2-fold acceleration ($l=2$) leads to a modest drop in metrics, with $n^{\text{cont}}=28$ achieving results comparable to the baseline. However, a 4-fold acceleration ($l=4$) significantly degrades quantitative performance. Despite this, qualitative analysis confirms that motion artifacts, such as tremors, are effectively mitigated, and visual quality remains acceptable. Interestingly for larger $l$ wider context improves the performance.
User study
To compare RDLA with our original method we conducted the user study of our DSG rolling method and RDLA with $l=2$, $n^{\text{cont}}=28$ in the same settings as in Sec. 5.3. The distribution of RDLA user study results is shown in Fig. 2 (Right). RDLA approach is only slightly inferior to our original method, which is consistent with the results of quantitative test.
Summary of acceleration findings
Our rolling modification of DSG achieves 10 FPS generation speed on a single Nvidia A40 GPU. With a ladder step of $l=1$ and $T_{r}=100$, our model achieves 70 FPS. Since it has lower latency while maintaining the same throughput as the DSG method, our approach is well-suited for an interaction with streaming audio. With $l=2$, the model achieves 120 FPS, further increasing real-time performance. Our experimental results demonstrate that RDLA can significantly enhance inference efficiency while maintaining high visual fidelity. By reducing denoising steps and employing ladder-based acceleration, our method achieves up to a 2 speedup. While minor metric degradation occurs at higher acceleration rates, RDLA remains a viable approach for real-time gesture synthesis applications.
6 Ablation Study
In our ablation study (see details in Supplemental Material), we analyzed key components influencing the Accelerated Rolling Diffusion framework’s performance. First, we evaluated the necessity of minimal noise in context frames. Results showed that incorporating minimal noise ($\sigma^{2}_{1}=0.00004$) into context frames significantly improved model robustness and prevented overfitting, thus enhancing generalization. Next, we assessed loss weighting strategies, comparing clamped-SNR [5] weighting versus uniform weighting. Uniform weighting ($a(t)=1$ in Eq. 5) provided a simpler, more stable training process without compromising performance, confirming its suitability for efficient sequential generation. Additionally, we conducted ablations on essential components of RDLA: OFS, inertial loss regularization, and progressive fine-tuning. Omitting OFS notably degraded gesture smoothness, leading to visible motion discontinuities. Removing inertial loss increased motion jitter, underscoring its importance for coherent motion. Progressive fine-tuning was crucial, as models without it exhibited significant quality and diversity loss. This strategy enabled effective adaptation to increased ladder step sizes, preserving performance.
7 Conclusion
We introduced Accelerated Rolling Diffusion, a novel framework enabling real-time, high-quality co-speech gesture generation through structured noise scheduling and Rolling Diffusion Ladder Acceleration (RDLA). Our method significantly enhances sampling efficiency, achieving up to a 2× speedup without compromising visual fidelity or temporal coherence. Extensive experiments on ZEGGS and BEAT benchmarks as well as user study confirm our framework’s generalizability and superior performance across various diffusion-based models. Thus, our approach represents a robust and efficient solution for streaming gesture synthesis, promising broad applicability in interactive embodied AI systems.
Supplementary Material
Appendix A RDLA: On-the-Fly Smoothing
Generalized version of On-the-Fly Smoothing (OFS). For $k\in\{0,1,\ldots,l/2\}$:
where $d_{c}=\cos{(\hat{x}_{n^{\text{cont}}+2k-2},\hat{x}_{n^{\text{cont}}+2k})}$.
Appendix B Experiments results: Accelerated inference via fewer diffusion steps
Results for reduced sampling steps on ZEGGS dataset using MSE metrics are in Tab. 5
Method, $T_{r}$ $\text{MSE}_{s}\downarrow$ $\text{MSE}_{k}\downarrow$
$\text{DSG}_{\text{DDPM}}$, $1000$ 23.42 143.13
$\text{Ours}_{\text{DDPM}}$, $1000$ 24.89 121.76
$\text{Ours}_{\text{DDPM}}$, $500$ 24.48 143.96
$\text{Ours}_{\text{DDIM}}$, $500$ 24.27 137.17
$\text{DSG}_{\text{DDPM}}$, $100$ 23.24 141.05
$\text{DSG}_{\text{DDIM}}$, $100$ 23.08 138.26
$\text{Ours}_{\text{DDPM}}$, $100$ 24.64 137.17
$\text{Ours}_{\text{DDIM}}$, $100$ 24.52 135.81
Appendix C RDLA experiments results
Results of RDLA across different ladder steps and context sizes on ZEGGS dataset using MSE metrics are in Tab. 6.
@@ Appendix C RDLA experiments results
Appendix D Ablation Study Details
@@
D.1 Noise in context frames
@@
Results in Tab. 7 and Tab. 8 indicate that completely removing noise from context frames leads to overfitting, because it required to adjust regularization parameters (the dropout rate increased from $0.2$ to $0.3$, the weight decay increased from $0.01$ to $0.1$) to achieve an acceptable result. But even with this adjustment, performance does not surpass the default setting where context frames contain minimal noise. This suggests that slight corruption of context frames during training improves robustness and generalization.
@@
D.2 Weighting in loss function
@@
Results in Tab. 9 and Tab. 10 show that clamped-SNR weighting does not improve performance over uniform weighting $a(t_{n})=1,\forall n$. Instead, uniform weighting consistently yields superior results. These findings contradict the hypothesis that assigning lower weight to higher-noise frames improves training stability. This suggests that equal importance across frames facilitates robust learning for sequential generation.
@@
D.3 RDLA ablations
To thoroughly evaluate the effectiveness of RDLA, we conduct an ablation study isolating key design components (Tab. 11 and Tab. 12). In all experiments $T=100$. One challenge introduced by RDLA is the potential for abrupt transitions between consecutive denoised blocks, leading to motion discontinuities. To mitigate this, we employ an On-the-Fly Smoothing (OFS) mechanism. As shown in Tab. 11 and Tab. 12, removing OFS results in a $9\%$ increase in $\text{FD}_{g}$, indicating a degradation in motion smoothness. This confirms that OFS effectively reduces visible discontinuities and enhances perceptual quality.
RDLA’s inertial loss function $L_{\text{inert}}$ is designed to regularize frame transitions during training. Ablation results in Tab. 11 and Tab. 12 show that excluding $L_{\text{inert}}$ increases $\text{FD}_{k}$ by $6\%$, indicating a loss in motion coherence. This suggests that incorporating an explicit regularization term improves the stability of frame transitions and reduces unwanted jitter in generated gestures.
Given that RDLA introduces structured noise scheduling with a variable step size, we adopt a progressive fine-tuning approach. Initially, the model is trained with a single-frame denoising schedule ($l=1$), followed by a gradual increase in $l$ (e.g., 2, 4, …). This iterative process enables the model to adapt to larger step sizes while minimizing performance degradation. To evaluate the effectiveness of this strategy, we compare models trained from scratch at $l=2$ against those fine-tuned from $l=1$. The results in Tab. 11 and Tab. 12 reveal that direct training at $l=2$ leads to much higher $\text{FD}_{g}$ and $\text{FD}_{k}$ values and reduced $\text{Div}_{g}$ and $\text{Div}_{k}$, underscoring the advantages of progressive adaptation. Moreover, bypassing training at larger step sizes results in a significant decline in performance, emphasizing the critical role of fine-tuning in preserving fidelity.
@@ D.3 RDLA ablations
Table 11: RDLA ablation study: impact of On-the-Fly Smoothing (OFS), Inertial loss ($L_{\text{inert}}$), fine-tuning from $l=1$ and bypassing training for $l=2$, $n^{\text{cont}}=28$. ’--’ indicates non-usage.
@@
-- Training 28.16 105.69
Table 12: RDLA ablation study using MSE metrics: impact of On-the-Fly Smoothing (OFS), Inertial loss ($L_{\text{inert}}$), fine-tuning from $l=1$ and bypassing training for $l=2$, $n^{\text{cont}}=28$. ’--’ indicates non-usage.

### v3

@@ Results
Participants were shown pairs of 15-second videos, each synchronized with the same audio but generated using different models. Both videos were displayed simultaneously, with their positions randomized in each trial to eliminate possible bias towards one side. The participants were asked to compare the two animations and rate them based on style consistency, naturalness and fluidity of animations, audio-animation synchronization, presence of technical issues (such as the gluing effect). The rating could take values $\{-2,-1,0,1,2\}$ where 2-2 indicates a strong preference for the baseline, 1-1 – slight baseline preference, 0 means no noticeable differences, $1$ indicates a slight preference for our model and $2$ indicates a strong preference for our model. To ensure the reliability of our evaluators, we included some videos multiple times and verified that the assessors provided consistent ratings before including their responses in the final analysis. The study was carried out on 60 pairs of video (10 per style in six different styles). Twenty-two professional assessors trained to work with video data were recruited to participate.
@@
The distribution of the user study results is shown in fig. 4 (Left). Our rolling modification of DSG significantly outperforms the original DSG, aligning well with the quantitative evaluation results; the preference is statistically significant (two-sided Wilcoxon signed-rank, $p<0.05$). To compare RDLA with our original method, we conducted a user study between DSG rolling and RDLA with $l=2$. The distribution of the RDLA user study results is shown in fig. 4 (Right). The RDLA approach is only slightly inferior to our original method, which is consistent with the quantitative findings.
5 Ablation Study
In our ablation study (see details in the Supplementary Material), we analyzed key components influencing the Accelerated Rolling Diffusion framework’s performance. First, we evaluated the necessity of minimal noise in context frames. Results showed that incorporating minimal noise ($\sigma^{2}_{1}=0.00004$) into context frames significantly improved model robustness and prevented overfitting, thus enhancing generalization. Next, we assessed loss weighting strategies, comparing clamped-SNR (barrault2024large) weighting versus uniform weighting. Uniform weighting ($a(t)=1$ in eq. 5) provided a simpler, more stable training process without compromising performance, confirming its suitability for efficient sequential generation. Additionally, we conducted ablations on the components of RDLA: inertial loss regularization (eq. 6), progressive fine-tuning, on-the-fly smoothing (OFS, not used in the reported results). Removing $L_{\text{RDLA}}(\theta)$ increased motion jitter, underscoring its importance for coherent motion. Progressive fine-tuning was crucial, as models without it exhibited significant quality and diversity loss. Omitting OFS only slightly degraded gesture smoothness. This strategy enabled effective adaptation to increased ladder step sizes, preserving performance.
6 Conclusion
We introduced Accelerated Rolling Diffusion, a framework for real-time, high-quality co-speech gesture generation. Experiments on ZEGGS and BEAT, plus user studies, demonstrate strong generalizability and consistent gains across diffusion-based gesture models. Our method produces continuous, temporally coherent motion without post-processing. To boost efficiency, we proposed Rolling Diffusion Ladder Acceleration (RDLA), which uses structured noise scheduling to reduce sampling steps and achieve up to 4× speedup while maintaining realism. RDLA is a promising direction for efficient diffusion acceleration, with potential extensions to other temporal modalities such as text and video.
Supplementary Material
Appendix A Related Work
Long-sequence Motion Generation. Generating long gesture sequences is challenging due to variable context-driven lengths and memory constraints. Gestures are typically conditioned by text, music, or speech, each requiring tailored approaches. Text-based methods often generate segments individually, stitching them together using weakly supervised techniques with motion smoothness priors (mao2022weaklysupervisedactiontransitionlearning; NEURIPS2023_2d52879e), iterative refinement like TEACH (10044438), or diffusion-based methods (e.g., DoubleTake) for temporal consistency (shafir2023humanmotiondiffusiongenerative). Precise temporal control is provided by multi-track timelines (petrovich2024multitracktimelinecontroltextdriven) and blended positional encodings (barquero2024seamlesshumanmotioncomposition). Additionally, hierarchical models like MotionMamba (zhang2024motionmambaefficientlong) and MultiAct (Lee_Moon_Lee_2023), along with language-driven frameworks such as MotionGPT (NEURIPS2023_3fbf0c1e), enable seamless, coherent generation across extended sequences.
Dancing Motions Generation. Long-sequence dance synthesis poses challenges due to duration and style diversity. Methods like EDGE (tseng2022edgeeditabledancegeneration) align overlapping clips for smooth transitions, while ”You Never Stop Dancing” (NEURIPS2022_40bfe617) uses low-dimensional manifolds with RefineBank and TransitBank to maintain fluid, high-quality motion.
Co-speech Gesture Generation. Driven by the pursuit of natural and expressive human-computer interactions, co-speech gesture generation has rapidly advanced using deep generative techniques. Researchers have explored methods from GANs, VAEs, and VQ-VAEs to diffusion models enhanced with transformer attention. While GANs (habibie2021learningspeechdriven3dconversational) often face instability and mode collapse, diffusion models deliver robustness, high fidelity, and diverse outputs. For instance, DiffuseStyleGesture (yang2023diffusestylegesture) integrates cross-local and self-attention to synchronize varied gestures, and EMAGE (Liu_2024_CVPR) employs masked audio-conditioned modeling with VQ-VAEs for greater expressivity. TalkSHOW (yi2023generatingholistic3dhuman) separately generates facial, body, and hand motions for nuanced speech alignment, while DiffTED (hogue2024difftedoneshotaudiodrivente) uses TPS keypoints in a diffusion pipeline to improve coherence. Lastly, Audio to Photoreal Embodiment (ng2024audiophotorealembodimentsynthesizing) combines diffusion with vector quantization to create realistic conversational avatars.
Long-sequence Co-Speech Gestures Generation Generating coherent long gesture sequences is challenging due to potential discontinuities and temporal inconsistencies. A common approach is to generate short chunks separately and then stitch overlapping clips together seamlessly (10.1145/3592458; Liu_2024_CVPR; yang2023diffusestylegesture; zhu2023tamingdiffusionmodelsaudiodriven; zhang2024speechdrivenpersonalizedgesturesynthetics). For example, FreeTalker (yang2024freetalkercontrollablespeechtextdriven) employs the DoubleTake blending technique to ensure smooth transitions, while autoregressive methods condition each new gesture on preceding outputs to preserve temporal coherence. Additionally, frameworks like DiffSHEG (chen2024diffsheg) use outpainting strategies to extend sequences incrementally, and DiffTED (hogue2024difftedoneshotaudiodrivente) generates coherent TPS keypoint sequences that bridge gesture synthesis with realistic video rendering.
Diffusion Models have gained popularity for generative tasks, particularly in image and video synthesis. Early work introduced denoising score matching (song2020generativemodelingestimatinggradients) and later evolved into denoising diffusion probabilistic models (DDPM) (ho2020denoising). Their success in high-quality image synthesis, seen in models like Imagen (saharia2022photorealistictexttoimagediffusionmodels) and Stable Diffusion (rombach2022highresolutionimagesynthesislatent), has been extended to video generation (ho2022videodiffusionmodels; bartal2024lumierespacetimediffusionmodel; ho2022imagenvideohighdefinition; blattmann2023stablevideodiffusionscaling). Given their temporal consistency and ability to model complex distributions, diffusion models are also being explored for co-speech gesture generation.
Rolling Diffusion Models (ruhe2024rollingdiffusionmodels) extend traditional diffusion-based generation to sequential data, enabling autoregressive synthesis by iteratively generating and conditioning on previous outputs. This approach enhances temporal consistency and long-range dependencies. Recent works (NEURIPS2024_2aee1c41; kim2024fifodiffusiongeneratinginfinitevideos; guo2025dynamical) continue to refine this concept making diffusion models more effective for sequential data generation by enhancing temporal consistency, enabling infinite-length generation, and improving temporal dynamics modeling.
Appendix B Rolling Diffusion Ladder Acceleration
On-the-Fly Smoothing
Our experiments revealed that although the RDLA approach significantly accelerates sampling, it may introduce motion artifacts in certain cases. Specifically, inconsistencies between denoised frame blocks may lead to tremors in motion, degrading both quantitative metrics and visual quality. To mitigate this, we introduce an On-the-Fly Smoothing (OFS) procedure, which refines transitions between consecutive denoised blocks. The core idea is to smooth the transition between the last frame of the previous block and the first frame of the newly denoised block. For $k\in\{0,1,\ldots,l/2\}$:
where $d_{c}=\cos{(\hat{x}_{n^{\text{cont}}+2k-2},\hat{x}_{n^{\text{cont}}+2k})}$, while $\tau$ is a predefined threshold controlling the degree of smoothing. The underlying intuition is that abrupt motion variations are most pronounced when successive frames exhibit high similarity, leading to a visual trembling effect. By averaging adjacent frames when necessary, OFS effectively suppresses these tremors without sacrificing the temporal integrity of motion sequences.
Appendix C Rolling Diffusion Ladder Acceleration Results
Rolling Diffusion Ladder Acceleration Results
@@ Appendix C Rolling Diffusion Ladder Acceleration Results
To evaluate the effectiveness of Rolling Diffusion Ladder Acceleration (RDLA) in improving inference efficiency, we conducted a series of experiments on the ZEGGS dataset (ghorbani2023zeroeggs). This dataset was selected due to its high fidelity, as it is obtained through motion capture, minimizing noise and annotation errors. To ensure a fair comparison, we applied RDLA to the DiffuseStyleGesture (DSG) model (yang2023diffusestylegesture), which has demonstrated state-of-the-art performance among diffusion-based methods on this benchmark. Following the evaluation protocol in (yang2023diffusestylegesture), we tested the performance of models on the ZEGGS validation set, which includes six distinct styles: Happy, Sad, Neutral, Old, Angry, and Relaxed. This results in a total of 36 audio samples for evaluation. In addition to the metrics presented in the main paper, we also report static and kinetic mean squared error (MSE) metrics here.
@@
Accelerated inference via fewer diffusion steps.
A straightforward method for accelerating the sampling process is reducing the number of denoising steps per frame. In our framework, each frame undergoes a predefined number of denoising steps $s$ before the rolling window shifts forward in time. We experimented with reducing this number $s_{r}$ from $s$ to 1 leading to a total inference step count of $T_{r}=s_{r}\cdot N$, where $N$ is the total number of frames in sliding window.
We compared the performance of our accelerated approach against the original DSG method under the same total denoising steps $T_{r}$. Additionally, we evaluated both DDPM (ho2020denoising) and DDIM (song2021denoising) sampling strategies, as DDIM is known to outperform DDPM when the number of steps gets smaller. Experiments were conducted with $N=100$, $n^{\text{cont}}=8$.
@@
Method, $T_{r}$ $\text{MSE}_{s}\downarrow$ $\text{MSE}_{k}\downarrow$
$\text{DSG}_{\text{DDPM}}$, $1000$ 23.42 143.13
$\text{Ours}_{\text{DDPM}}$, $1000$ 24.89 121.76
$\text{Ours}_{\text{DDIM}}$, $1000$ 25.02 132.54
$\text{Ours}_{\text{DDPM}}$, $500$ 24.48 143.96
$\text{Ours}_{\text{DDIM}}$, $500$ 24.27 137.17
$\text{DSG}_{\text{DDPM}}$, $100$ 23.24 141.05
$\text{DSG}_{\text{DDIM}}$, $100$ 23.08 138.26
$\text{Ours}_{\text{DDPM}}$, $100$ 24.64 137.17
$\text{Ours}_{\text{DDIM}}$, $100$ 24.52 135.81
As shown in table 4 and table 5, reducing the number of denoising steps generally maintains and even improves performance in both Fréchet distance (FD) and diversity (Div) metrics, with our approach consistently outperforming the baseline DSG method. Notably, while DDIM achieves competitive results, DDPM remains superior in most cases. These findings suggest that our method can effectively operate with a reduced number of denoising steps (e.g. $T_{r}=100$) without significant degradation in quality.
Temporal acceleration experiments.
Beyond reducing the number of sampling steps, we leveraged RDLA to accelerate inference in the temporal dimension. By constructing a denoising ladder with step size $l$, we simultaneously denoised $l$ frames at each iteration, achieving an $l$-fold acceleration. Our experiments applied RDLA to DSG, using $N+n^{cont}=108$, ladder step sizes $l\in\{2,4\}$, and various context sizes $n^{\text{cont}}\in\{8,18,28,38\}$. To maximize speedup, we coupled RDLA with reduced denoising steps, using $T_{r}\in\{100,90,80,70\}$, corresponding to different $n^{\text{cont}}$ values.
@@
Results in table 6 and table 7 indicate that a 2-fold acceleration ($l=2$) leads to a modest drop in metrics, with $n^{\text{cont}}=28$ achieving results comparable to the baseline. However, a 4-fold acceleration ($l=4$) degrades quantitative performance. Despite this, qualitative analysis confirms that motion artifacts are effectively mitigated, and visual quality remains acceptable. Interestingly for larger $l$ wider context improves the performance.
Appendix D Ablation Study
@@
Noise in context frames
@@
Results in table 8 and table 9 indicate that completely removing noise from context frames leads to overfitting, because it required to adjust regularization parameters (the dropout rate increased from $0.2$ to $0.3$, the weight decay increased from $0.01$ to $0.1$) to achieve an acceptable result. But even with this adjustment, performance does not surpass the default setting where context frames contain minimal noise. This suggests that slight corruption of context frames during training improves robustness and generalization.
@@
Weighting in loss function
@@
Results in table 10 and table 11 show that clamped-SNR weighting does not improve performance over uniform weighting $a(t_{n})=1,\forall n$. Instead, uniform weighting consistently yields superior results. These findings contradict the hypothesis that assigning lower weight to higher-noise frames improves training stability. This suggests that equal importance across frames facilitates robust learning for sequential generation.
@@
RDLA ablations
To thoroughly evaluate the effectiveness of RDLA, we conduct an ablation study isolating key design components (table 12 and table 13). In all experiments $T=100$. One challenge introduced by RDLA is the potential for abrupt transitions between consecutive denoised blocks, leading to motion discontinuities. To mitigate this, we employ an On-the-Fly Smoothing (OFS) mechanism. As shown in table 12 and table 13, removing OFS results in a slight degradation in motion smoothness.
RDLA’s inertial loss function $L_{\text{inert}}$ is designed to regularize frame transitions during training. Ablation results in table 12 and table 13 show that excluding $L_{\text{inert}}$ increases $\text{FD}_{k}$ by $6\%$, indicating a loss in motion coherence. This suggests that incorporating an explicit regularization term improves the stability of frame transitions and reduces unwanted jitter in generated gestures.
Given that RDLA introduces structured noise scheduling with a variable step size, we adopt a progressive fine-tuning approach. Initially, the model is trained with a single-frame denoising schedule ($l=1$), followed by a gradual increase in $l$ (e.g., 2, 4, …). This iterative process enables the model to adapt to larger step sizes while minimizing performance degradation. To evaluate the effectiveness of this strategy, we compare models trained from scratch at $l=2$ against those fine-tuned from $l=1$. The results in table 12 and table 13 reveal that direct training at $l=2$ leads to much higher $\text{FD}_{g}$ and $\text{FD}_{k}$ values and reduced $\text{Div}_{g}$ and $\text{Div}_{k}$, underscoring the advantages of progressive adaptation. Moreover, bypassing training at larger step sizes results in a significant decline in performance, emphasizing the critical role of fine-tuning in preserving fidelity.
@@
Table 12: RDLA ablation study: impact of On-the-Fly Smoothing (OFS), Inertial loss ($L_{\text{RDLA}}(\theta)$), fine-tuning from $l=1$ and bypassing training for $l=2$, $n^{\text{cont}}=28$. ’-’ indicates non-usage.
@@
- Training 28.16 505.69
Table 13: RDLA ablation study using MSE metrics: impact of On-the-Fly Smoothing (OFS), Inertial loss ($L_{\text{RDLA}}(\theta)$), fine-tuning from $l=1$ and bypassing training for $l=2$, $n^{\text{cont}}=28$. ’-’ indicates non-usage.
Appendix E Qualitative Comparisons
Figure 5: Qualitative Comparison on ZEGGS Dataset. Columns (a–c) show DiffuseStyleGesture, Taming, and PersonaGestor; the baseline is on the bottom row and our version on the top. Our method produces natural, dynamic gestures that are tightly aligned with the speech content. Notably, it uses an energetic hand movement to point toward the interlocutor on the word “you,” effectively emphasizing the rhetorical intent of the phrase “can you think for a second, that that’s a good idea,” while the baselines largely lack emphasis and gesture variation.
Figure 6: Qualitative Comparison on ZEGGS Dataset. Columns (a–c) show DiffuseStyleGesture, Taming, and PersonaGestor; the baseline is on the bottom row and our version on the top. Our approach generates rhythmically aligned and semantically expressive gestures. In particular, it uses distinct hand movements to accompany and emphasize the repeated structure in “can eat there for breakfast, can eat there for lunch,” effectively conveying a sense of enumeration and emphasis, while the baselines remain mostly static or repetitive.